### v1

# Constructing new probability distributions on the unit interval

Abstract
In this paper, we introduce a novel method for constructing probability distributions on the unit interval by exploiting the non-injective transformation defined by the ratio of two positive random variables, $X$ and $Y$. For simplicity and tractability, we focus on independent random variables $X$ and $Y$ following gamma distributions. We derive some results including laws, density functions, quantile function and closed-form expressions for the moments, and discuss maximum likelihood estimation.
Keywords. Gamma distribution, Unit probabilistic model, Maximum likelihood.
Mathematics Subject Classification (2010). MSC 60E05 $\cdot$ MSC 62Exx $\cdot$ MSC 62Fxx.
1 Introduction
Probability models with suppport on the unit interval are usefull in many fields, including econometrics (Cribari-Neto and Zeileis,, 2010), natural sciences Geissinger et al., (2022), and reliability Cross et al., (2006), among others. Traditionally, the beta (Ferrari and Cribari‐Neto,, 2004) and Kumaraswamy (Kumaraswamy,, 1980) distributions have been the workhorses for modeling such data, due to their flexibility and tractability. More recently, alternative unit‐interval families have emerged, see, for example, Vila et al., 2024c ; Vila et al., 2024b , to address features like bimodality.
The beta distribution is constructed by the ratio of two positive random variables, that is, when $X,Y\sim\mathrm{Gamma}(\alpha_{i},\lambda)$ are independent and identically scaled, then
$\frac{X}{X+Y}\sim\mathrm{Beta}(\alpha_{1},\alpha_{2}),$
yielding the familiar beta model. In this work, we introduce two new families of unit‐interval distributions, defined by non‐injective transformations of the ratio $X/(X+Y)$. Writing
[…]
The remainder of this paper is organized as follows. In Section 2, we introduce the new unit‐interval families via the non‐injective transformations $S_{r}$ and $T_{r}$ of the gamma ratio. In Section 3, we derive their distributional laws in terms of the underlying $X/(X+Y)$ density. In Section 4, we obtain closed‐form expressions for the density and cumulative distribution functions of $W$ and $Z$. In Section 5, we present the quantile function. In Section 6, we develop closed‐form formulae for selected raw moments under special parameter regimes. In Section 7, we formulate maximum likelihood estimation, derive the likelihood equations, and discuss numerical solution strategies. Finally, in Section 8, we provide some concluding remarks.
2 The unit probabilistic models
Let $X$ and $Y$ be two (absolutely) continuous random variables, with positive support, both defined in the same probability space $(\Omega,\mathscr{B}(\Omega),\mathbb{P})$. Here, $\mathscr{B}(\Omega)$ denotes the Borel $\sigma$-algebra on $\Omega$ and $\mathbb{P}:\mathscr{B}(\Omega)\to[0,1]$ is a probability measure. In this work we define news random variables as the composition of a surjective transformation $\varphi_{r}:(0,1)\to(0,1)$ with the ratio $X/(X+Y)$, that is,
$\varphi_{r}\circ\left({X\over X+Y}\right)=\varphi_{r}\left({X\over X+Y}\right):\Omega\to(0,1),$ (1)
so that the addition of the parameter $r>0$ generates flexibility in the proposed models.
Note that the complexity of the probabilistic models defined in (1) depends on the choices of variables $X$ and $Y$ and on its correlation structure, which are innumerable. As an illustration, reference (Colin and Ossikovski,, 2025) explores various probabilistic aspects of model (1) under the conditions $\varphi_{r}(x)=[1-4x(1-x)]^{r}$ and $r=1/2$, assuming that the variables $X$ and $Y$ are correlated Gamma-distributed with the same shape and scale parameters, and a linear correlation coefficient $\rho$. For simplicity of presentation, from now on we choose $X$ and $Y$ independent (not necessarily identically distributed) with gamma distributions, denoted by $X\sim\text{Gamma}(\alpha_{1},\lambda_{1})$, $\alpha_{1},\lambda_{1}>0$, and $Y\sim\text{Gamma}(\alpha_{2},\lambda_{2})$, $\alpha_{2},\lambda_{2}>0$, respectively. That is, the density functions of $X$ and $Y$ are given by
$f_{X}(x)={\lambda_{1}^{\alpha_{1}}\over\Gamma(\alpha_{1})}\,x^{\alpha_{1}-1}\exp(-\lambda_{1}x),\quad f_{Y}(y)={\lambda_{2}^{\alpha_{2}}\over\Gamma(\alpha_{2})}\,y^{\alpha_{2}-1}\exp(-\lambda_{2}y),\quad x,y>0,$
[…]
${X\over X+Y}\sim\text{Beta}(\alpha_{1},\alpha_{2}).$
[…]
$S_{r}(x)\equiv\varphi_{r}(x)=[1-4x(1-x)]^{r},\quad T_{r}(x)\equiv\varphi_{r}(x)=1-[4x(1-x)]^{r},\quad 0<x<1;\ r>0.$ (2)
Note that both transformations above have a zero at the point $x=1/2$ and assume the value of 111 at the points $x=0$ and $x=1$, see Figure 1.
Figure 1: Plots of the transformations $S_{r}$ and $T_{r}$ for different choices of the parameter $r$.
Using the transformations $S_{r}$ and $T_{r}$ in (2), simple algebraic manipulations show that, the new random variables $W$ and $Z$, defined as
$W\equiv S_{r}\left({X\over X+Y}\right)=\left({Y-X\over X+Y}\right)^{2r},$ (3)
$Z\equiv T_{r}\left({X\over X+Y}\right)=1-{(XY)^{r}\over\displaystyle\left[{1\over 2}(X+Y)\right]^{2r}},$ (4)
respectively, have support in the unit interval.
Remark 2.1.
Note that when $r=1$, we get $W=Z$.
Remark 2.2.
The value of $r=1/2$ in formulas (3) and (4) is special, because for this value the variables $W$ and $Z$ take the forms
$W={|X-Y|\over X+Y}\quad\text{and}\quad Z=1-{\sqrt{XY}\over{1\over 2}(X+Y)},$ (5)
which coincide, respectively, with the Gini coefficient (Gini,, 1936) and Atkinson index (Atkinson,, 1970) estimators formed by the random sample $(X,Y)$ of size $n=2$.
3 Deriving the laws of $W$ and $Z$
Note that, for any Borel set $B$ on $\mathbb{R}$, we have
$\mathbb{P}\left(W\in B\right)=\mathbb{P}\left(S_{r}\left({X\over X+Y}\right)\in B\right)$ =($=\mathbb{P}\left(\left\{\omega\in\Omega:{X(\omega)\over X(\omega)+Y(\omega)}\in S_{r}^{-1}(B)\right\}\right)$
$=\int_{S_{r}^{-1}(B)}f_{X\over X+Y}(u){\rm d}u,$ (6)
where $S_{r}^{-1}(B)=\{x\in\mathbb{R}:S_{r}(x)\in B\}$ is the inverse image of set $B$ under the function $S_{r}$. Since the Borel $\sigma$-algebra on $\mathbb{R}$ is generated by open or closed sets, without loss of generality, assume that $B=(a,b)$, with $a<b$. Hence,
$S_{r}^{-1}(B)=\begin{cases}\left({1-b_{*}^{{}^{1\over 2r}}\over 2},{1-a_{*}^{{}^{1\over 2r}}\over 2}\right)\bigcup\left({1+a_{*}^{{}^{1\over 2r}}\over 2},{1+b_{*}^{{}^{1\over 2r}}\over 2}\right),&B\cap(0,1)\neq\emptyset,\\[8.5359pt] \emptyset,&B\cap(0,1)=\emptyset,\end{cases}$
where $a_{*}=a\vee 0=\max\{a,0\}$ and $b_{*}=b\wedge 1=\min\{b,1\}$. Therefore, by using (3), the law of $W$ can be written as
$\mathbb{P}\left(W\in B\right)=\begin{cases}\displaystyle\int_{{\left({1-b_{*}^{{}^{1\over 2r}}\over 2},{1-a_{*}^{{}^{1\over 2r}}\over 2}\right)\bigcup\left({1+a_{*}^{{}^{1\over 2r}}\over 2},{1+b_{*}^{{}^{1\over 2r}}\over 2}\right)}}f_{X\over X+Y}(u){\rm d}u,&B\cap(0,1)\neq\emptyset,\\[8.5359pt] 0,&B\cap(0,1)=\emptyset.\end{cases}$ (7)
Similarly, we can show that the law of $Z$ can be expressed as
$\mathbb{P}\left(Z\in B\right)=\begin{cases}\displaystyle\int_{{\left({1-\sqrt{1-(1-b_{*})^{1\over r}}\over 2},{1-\sqrt{1-(1-a_{*})^{1\over r}}\over 2}\right)\bigcup\left({1+\sqrt{1-(1-a_{*})^{1\over r}}\over 2},{1+\sqrt{1-(1-b_{*})^{1\over r}}\over 2}\right)}}f_{X\over X+Y}(u){\rm d}u,&B\cap(0,1)\neq\emptyset,\\[8.5359pt] 0,&B\cap(0,1)=\emptyset.\end{cases}$ (8)
By (7) and (8), we note that knowledge of the laws of $W$ and $Z$ relies on knowledge of density of the independent ratio $X/(X+Y)$, with $X\sim\text{Gamma}(\alpha_{1},\lambda_{1})$ and $Y\sim\text{Gamma}(\alpha_{2},\lambda_{2})$, which is well-known in literature (see Example 6 of reference Vila et al., 2024a, ):
$f_{X\over X+Y}(v)={1\over\lambda_{1}^{\alpha_{2}}\lambda_{2}^{\alpha_{1}}\text{B}(\alpha_{1},\alpha_{2})}\,\dfrac{v^{\alpha_{1}-1}(1-v)^{\alpha_{2}-1}}{\displaystyle\left[\left({1\over\lambda_{2}}-{1\over\lambda_{1}}\right)v+{1\over\lambda_{1}}\right]^{\alpha_{1}+\alpha_{2}}},\quad 0<v<1,$ (9)
where $\text{B}(z_{1},z_{2})=\int_{0}^{1}t^{z_{1}-1}(1-t)^{z_{2}-1}{\rm d}t$ is the (complete) beta function.
4 Distribution and density functions of $W$ and $Z$
By taking $a\to-\infty$ and $b=w\in(0,1)$ in (7), we have $a_{*}=0$ and $b_{*}=w$. Then the cumulative distribution function (CDF) of $W$ can be written as
$F_{W}(w)=\int_{1-w_{r}}^{w_{r}}f_{X\over X+Y}(u){\rm d}u,\quad 0<w<1,$ (10)
where we have adopted the notation
$w_{r}\equiv{1+w^{\!{}^{1\over 2r}}\over 2}.$ (11)
[…]
$f_{W}(w)={w^{{1\over 2r}-1}\over 4r}\left[f_{X\over X+Y}\left(w_{r}\right)+f_{X\over X+Y}\left(1-w_{r}\right)\right].$ (12)
Analogously, by taking $a\to-\infty$ and $b=z\in(0,1)$ in (8), we have $a_{*}=0$ and $b_{*}=z$. Then the CDF of $Z$ can be written as
$F_{Z}(z)=\int_{1-z_{r}}^{z_{r}}f_{X\over X+Y}(u){\rm d}u,\quad 0<z<1,$ (13)
where we are using the notation
$z_{r}\equiv{1+\sqrt{1-(1-z)^{1\over r}}\over 2}.$ (14)
[…]
$f_{Z}(z)={(1-z)^{{1\over r}-1}\over 4r\sqrt{1-(1-z)^{1\over r}}}\left[f_{X\over X+Y}\left(z_{r}\right)+f_{X\over X+Y}\left(1-z_{r}\right)\right].$ (15)
Plugging (9) into formulas (12) and (15), we get (for $0<w,z<1$)
$f_{W}(w)={1\over 4r\lambda_{1}^{\alpha_{2}}\lambda_{2}^{\alpha_{1}}\text{B}(\alpha_{1},\alpha_{2})}\,w^{{1\over 2r}-1}G(w_{r})$ (16)
and
$f_{Z}(z)$ $={1\over 4r\lambda_{1}^{\alpha_{2}}\lambda_{2}^{\alpha_{1}}\text{B}(\alpha_{1},\alpha_{2})}\,{(1-z)^{{1\over r}-1}G(z_{r})\over\sqrt{1-(1-z)^{1\over r}}},$ (17)
respectively, where
$G(x)\equiv\dfrac{\displaystyle x^{\alpha_{1}-1}\left(1-x\right)^{\alpha_{2}-1}}{\displaystyle\left[\left({1\over\lambda_{2}}-{1\over\lambda_{1}}\right)x+{1\over\lambda_{1}}\right]^{\alpha_{1}+\alpha_{2}}}+\dfrac{\displaystyle\left(1-x\right)^{\alpha_{1}-1}x^{\alpha_{2}-1}}{\displaystyle\left[\left({1\over\lambda_{2}}-{1\over\lambda_{1}}\right)(1-x)+{1\over\lambda_{1}}\right]^{\alpha_{1}+\alpha_{2}}}.$ (18)
Figures 2-7 display a range of shapes for the densities of $W$ and $Z$.
Figure 2: Plots of the PDFs $f_{W}$ (left) and $f_{Z}$ (right) with varying parameters $\alpha_{1},\alpha_{2},\lambda_{1}$ and $\lambda_{2}$, with $r=0.3$.
Figure 3: Plots of the PDFs $f_{W}$ (left) and $f_{Z}$ (right) with varying parameters $\alpha_{1},\alpha_{2},\lambda_{1}$ and $\lambda_{2}$, with $r=0.4$.
Figure 4: Plots of the PDFs $f_{W}$ (left) and $f_{Z}$ (right) with varying parameters $\alpha_{1},\alpha_{2},\lambda_{1}$ and $\lambda_{2}$, with $r=0.5$.
Figure 5: Plots of the PDFs $f_{W}$ (left) and $f_{Z}$ (right) with varying parameters $\alpha_{1},\alpha_{2},\lambda_{1}$ and $\lambda_{2}$, with $r=1$.
Figure 6: Plots of the PDFs $f_{W}$ (left) and $f_{Z}$ (right) with varying parameters $\alpha_{1},\alpha_{2},\lambda_{1}$ and $\lambda_{2}$, with $r=2$.
Figure 7: Plots of the PDFs $f_{W}$ (left) and $f_{Z}$ (right) with varying parameters $\alpha_{1},\alpha_{2},\lambda_{1}$ and $\lambda_{2}$, with $r=3$.
Remark 4.1.
In the special case $\lambda_{1}=\lambda_{2}$, we get
$f_{W}(w)={w^{{1\over 2r}-1}\over 4r\text{B}(\alpha_{1},\alpha_{2})}\left[{\displaystyle w_{r}^{\alpha_{1}-1}(1-w_{r})^{\alpha_{2}-1}}\right.+\left.{\displaystyle(1-w_{r})^{\alpha_{1}-1}w_{r}^{\alpha_{2}-1}}\right],\quad 0<w<1,$
and
$f_{Z}(z)$ $={(1-z)^{{1\over r}-1}\over 4r\text{B}(\alpha_{1},\alpha_{2})\sqrt{1-(1-z)^{1\over r}}}\left[{\displaystyle z_{r}^{\alpha_{1}-1}(1-z_{r})^{\alpha_{2}-1}}+{\displaystyle(1-z_{r})^{\alpha_{1}-1}z_{r}^{\alpha_{2}-1}}\right],\quad 0<z<1,$
where $w_{r}$ and $z_{r}$ are as given in (11) and (14), respectively.
Remark 4.2.
It is worth noting that in the case $\lambda_{1}=\lambda_{2}\equiv\lambda$, the densities $f_{W}$ and $f_{Z}$ presented in Remark 4.1 do not depend on $\lambda$, consistent with the fact that $W$ and $Z$, given in (3) and (4), respectively, are derived from scale-invariant transformations. That is, for $X\sim\text{Gamma}(\alpha_{1},\lambda)$ and $Y\sim\text{Gamma}(\alpha_{2},\lambda)$ we have $\lambda X\sim\text{Gamma}(\alpha_{1},1)$ and $\lambda Y\sim\text{Gamma}(\alpha_{2},1)$, and consequently, the random variables $W$ and $Z$ satisfy the identities
$W=\left({Y-X\over X+Y}\right)^{2r}=\left({\lambda Y-\lambda X\over\lambda X+\lambda Y}\right)^{2r},$
$Z=1-{(XY)^{r}\over\displaystyle\left[{1\over 2}(X+Y)\right]^{2r}}=1-{(\lambda X\lambda Y)^{r}\over\displaystyle\left[{1\over 2}(\lambda X+\lambda Y)\right]^{2r}},$
whose distribution is unaffected by $\lambda$.
5 Quantiles
From (10), the $p$-quantile function $Q_{W}(p)$ of $W$, for given $p\in(0,1)$, satisfies the relation
$F_{W}(Q_{W}(p))=F_{X\over X+Y}\left({1+Q_{W}^{\!{}^{1\over 2r}}(p)\over 2}\right)-F_{X\over X+Y}\left({1-Q_{W}^{\!{}^{1\over 2r}}(p)\over 2}\right).$
Analogously, from (13), the $p$-quantile function $Q_{Z}(p)$ of $Z$, for given $p\in(0,1)$, satisfies
$F_{Z}(Q_{Z}(p))=F_{X\over X+Y}\left({1+\sqrt{1-[1-Q_{Z}(p)]^{1\over r}}\over 2}\right)-F_{X\over X+Y}\left({1-\sqrt{1-[1-Q_{Z}(p)]^{1\over r}}\over 2}\right).$
While $Q_{W}(p)$ and $Q_{Z}(p)$ can be theoretically derived from the above equations, their expressions lack simple closed forms, necessitating numerical methods to invert these equations for quantiles of $W$ and $Z$.
6 Closed-form expressions for the $n$th raw moments of $W$ and $Z$
It is clear that, by (16) and (17), the $n$th raw moments of $W$ and $Z$ can be calculated by the following formulas:
$\mathbb{E}(W^{n})$ $={1\over 4r\lambda_{1}^{\alpha_{2}}\lambda_{2}^{\alpha_{1}}\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{1}w^{n+{1\over 2r}-1}G(w_{r}){\rm d}w$
and
$\mathbb{E}(Z^{n})$ $={1\over 4r\lambda_{1}^{\alpha_{2}}\lambda_{2}^{\alpha_{1}}\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{1}{z^{n}(1-z)^{{1\over r}-1}G(z_{r})\over\sqrt{1-(1-z)^{1\over r}}}{\rm d}z,$
where $w_{r}$ and $z_{r}$ are as given in (11) and (14), respectively, and $G(\cdot)$ is as in (18). Note that the above integrals are well-defined due to the support of $W$ and $Z$ in (0,1)(0,1)( 0 , 1 ), but numerical methods are necessary for their evaluation due to analytical complexity.
The following Subsections 6.1, 6.2, 6.3 and 6.4 provide some analytical expressions for the raw moments of $W$ and $Z$ subject to parameter constraints.
6.1 The 111st raw moment of $W$: case $r=1/2$ and $1/3<\lambda_{2}/\lambda_{1}<2$
Under the conditions $r=1/2$ and $1/3<\lambda_{2}/\lambda_{1}<2$, in this part, we obtain (more precisely, in (25) and Remark 6.2) straightforward closed-form expressions for the 111st raw moment (expected value) of
$W=\frac{|X-Y|}{X+Y},$
where $X\sim\text{Gamma}(\alpha_{1},\lambda_{1})$ and $Y\sim\text{Gamma}(\alpha_{2},\lambda_{2})$ are independent.
Indeed, as the function $(0,\infty)\ni x\mapsto z\exp(-zx)$, $z>0$, is the density of an exponential distribution, we have
$\int_{0}^{\infty}\exp(-zx){\rm d}x={1\over z}.$ (19)
By making $z=X+Y$ and by using the fact that $X$ and $Y$ are independent, we get
$\mathbb{E}(W)=\mathbb{E}\left[\frac{|X-Y|}{X+Y}\right]$ $=\mathbb{E}\left[|X-Y|\int_{0}^{\infty}\exp\left\{-\left(X+Y\right)x\right\}{\rm d}x\right]$
$=\int_{0}^{\infty}\mathbb{E}\left[|X-Y|\exp\left\{-(X+Y)x\right\}\right]{\rm d}x,$ (20)
where in the last step, we applied the Tonelli’s theorem (Folland,, 1999) to justify the change in the order of integration.
On the other hand, note that
$\mathbb{E}\left[|X-Y|\exp\left\{-(X+Y)x\right\}\right]$ $=\int_{0}^{\infty}\left[\int_{u}^{\infty}(v-u)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$
$+\int_{0}^{\infty}\left[\int_{0}^{u}(u-v)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$
$=\int_{0}^{\infty}\left[\int_{0}^{\infty}(v-u)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$
$-\int_{0}^{\infty}\left[\int_{0}^{u}(v-u)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$
$+\int_{0}^{\infty}\left[\int_{0}^{u}(u-v)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u).$ (21)
Since
$\int_{0}^{\infty}\Bigg{[}$ $\int_{0}^{\infty}(v-u)\exp(-vx){\rm d}F_{Y}(v)\Bigg{]}\exp(-ux)=\int_{0}^{\infty}\exp(-ux){\rm d}F_{X}(u)\int_{0}^{\infty}v\exp(-vx){\rm d}F_{Y}(v)$
$-\int_{0}^{\infty}u\exp(-ux){\rm d}F_{X}(u)\int_{0}^{\infty}\exp(-vx){\rm d}F_{Y}(v)$
$=\mathbb{E}\left[\exp(-Xx)\right]\mathbb{E}\left[Y\exp(-Yx)\right]-\mathbb{E}\left[X\exp(-Xx)\right]\mathbb{E}\left[\exp(-Yx)\right]$
and
$-\int_{0}^{\infty}\left[\int_{0}^{u}(v-u)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$
$+\int_{0}^{\infty}\left[\int_{0}^{u}(u-v)\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$
$=2\int_{0}^{\infty}\left[\int_{0}^{u}\exp(-vx){\rm d}F_{Y}(v)\right]u\exp(-ux){\rm d}F_{X}(u)$
$-2\int_{0}^{\infty}\left[\int_{0}^{u}v\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u),$
the identity in (21) becomes
$\mathbb{E}\left[|X-Y|\exp\left\{-(X+Y)x\right\}\right]$ $=\mathbb{E}\left[\exp(-Xx)\right]\mathbb{E}\left[Y\exp(-Yx)\right]-\mathbb{E}\left[X\exp(-Xx)\right]\mathbb{E}\left[\exp(-Yx)\right]$
$+2\int_{0}^{\infty}\left[\int_{0}^{u}\exp(-vx){\rm d}F_{Y}(v)\right]u\exp(-ux){\rm d}F_{X}(u)$
$-2\int_{0}^{\infty}\left[\int_{0}^{u}v\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u),$ (22)
where $X\sim\text{Gamma}(\alpha_{1},\lambda_{1})$ and $Y\sim\text{Gamma}(\alpha_{2},\lambda_{2})$.
By using results of Propositions A.1 and A.2 in (22), it follows that
$\mathbb{E}\left[|X-Y|\exp\left\{-(X+Y)x\right\}\right]$ $={\alpha_{2}\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over(x+\lambda_{1})^{\alpha_{1}}(x+\lambda_{2})^{\alpha_{2}+1}}-{\alpha_{1}\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over(x+\lambda_{1})^{\alpha_{1}+1}(x+\lambda_{2})^{\alpha_{2}}}$
$+{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2}+1)}\,{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2}+1,1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}+1}}$
$-{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\,{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2},1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}$
$+{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\,{1\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}.$ (23)
By replacing (23) in (20), we get
$\mathbb{E}(W)$ $=\alpha_{2}\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\int_{0}^{\infty}{1\over(x+\lambda_{1})^{\alpha_{1}}(x+\lambda_{2})^{\alpha_{2}+1}}{\rm d}x-\alpha_{1}\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\int_{0}^{\infty}{1\over(x+\lambda_{1})^{\alpha_{1}+1}(x+\lambda_{2})^{\alpha_{2}}}{\rm d}x$
$+{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2}+1)}\int_{0}^{\infty}{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2}+1,1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}+1}}{\rm d}x$
$-{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{\infty}{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2},1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}{\rm d}x$
$+{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{\infty}{1\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}{\rm d}x.$ (24)
By using Proposition A.3, the expectation of $W$ in (24) can be written as
$\mathbb{E}(W)$ $={\alpha_{2}\lambda_{2}^{\alpha_{2}}\over\lambda_{1}^{\alpha_{2}}(\alpha_{1}+\alpha_{2})}\,_{2}F_{1}\left(\alpha_{2}+1,\alpha_{1}+\alpha_{2};\alpha_{1}+\alpha_{2}+1;1-{\lambda_{2}\over\lambda_{1}}\right)$
$-{\alpha_{1}\lambda_{2}^{\alpha_{2}}\over\lambda_{1}^{\alpha_{2}}(\alpha_{1}+\alpha_{2})}\,_{2}F_{1}\left(\alpha_{2},\alpha_{1}+\alpha_{2};\alpha_{1}+\alpha_{2}+1;1-{\lambda_{2}\over\lambda_{1}}\right)$
$+{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2}+1)}\int_{0}^{\infty}{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2}+1,1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}+1}}{\rm d}x$
$-{2\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\int_{0}^{\infty}{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2},1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}{\rm d}x$
$+{\lambda_{1}^{\alpha_{1}}\over 2^{\alpha_{1}+\alpha_{2}-1}\lambda_{2}^{\alpha_{1}}(\alpha_{1}+\alpha_{2})\text{B}(\alpha_{1},\alpha_{2})}\,_{2}F_{1}\left(\alpha_{1}+\alpha_{2},\alpha_{1}+\alpha_{2};\alpha_{1}+\alpha_{2}+1;1-{\lambda_{1}+\lambda_{2}\over 2\lambda_{2}}\right),$ (25)
provided $1/3<\lambda_{2}/\lambda_{1}<2$.
Remark 6.1.
It is noteworthy that, for $\lambda_{1}\neq\lambda_{2}$, the integrals in (25) lack closed-form representations in terms of standard special functions, but can be efficiently computed numerically. For the special case $\lambda_{1}=\lambda_{2}$, Remark 6.2 provides a straightforward closed-form expression for the expected value of $W$.
Remark 6.2.
Setting $\lambda_{1}=\lambda_{2}$ in (25) and using Proposition A.3, we obtain.
$\mathbb{E}(W)={\alpha_{2}-\alpha_{1}\over\alpha_{1}+\alpha_{2}}+{1\over 2^{\alpha_{1}+\alpha_{2}-1}(\alpha_{1}+\alpha_{2})}\\[5.69046pt] \times{1\over\text{B}(\alpha_{1},\alpha_{2})}\left[1+{\alpha_{1}+\alpha_{2}\over 2\alpha_{2}}\,_{2}F_{1}\left(\alpha_{1}+\alpha_{2}+1,1;\alpha_{2}+1;{1\over 2}\right)\right.-\left.\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2},1;\alpha_{2}+1;{1\over 2}\right)\right],$
because $\,{}_{2}F_{1}(a,b;c;0)=1$ and $\text{B}(\alpha_{1},\alpha_{2}+1)=\alpha_{2}\text{B}(\alpha_{1},\alpha_{2})/(\alpha_{1}+\alpha_{2})$. If furthermore, $\alpha_{1}=\alpha_{2}\equiv\alpha$, the above expression reduces to
$\mathbb{E}(W)={\Gamma(2\alpha)\over 2^{2\alpha}\alpha\Gamma^{2}(\alpha)}\left[1+\,_{2}F_{1}\left(2\alpha+1,1;\alpha+1;{1\over 2}\right)\right.-\left.\,{}_{2}F_{1}\left(2\alpha,1;\alpha+1;{1\over 2}\right)\right].$ (26)
By applying the identity (Wolfram Research, Inc.,, 2024):
$\,{}_{2}F_{1}\left(a,b;{a+b\over 2};{1\over 2}\right)=\sqrt{\pi}\Gamma\left({a+b\over 2}\right)\left[{1\over\Gamma({a+1\over 2})\Gamma({b\over 2})}+{1\over\Gamma({a\over 2})\Gamma({b+1\over 2})}\right],$
we have
$\,_{2}F_{1}\left(2\alpha+1,1;\alpha+1;{1\over 2}\right)=1+{\sqrt{\pi}\Gamma\left(\alpha+1\right)\over\Gamma({2\alpha+1\over 2})}.$ (27)
Furthermore, by using the identity (Wolfram Research, Inc.,, 2024):
$\,{}_{2}F_{1}\left(a,b;{a+b+1\over 2};{1\over 2}\right)={\sqrt{\pi}\Gamma({a+b+1\over 2})\over\Gamma({a+1\over 2})\Gamma({b+1\over 2})},$
we get
$\,_{2}F_{1}\left(2\alpha,1;\alpha+1;{1\over 2}\right)={\sqrt{\pi}\Gamma(\alpha+1)\over\Gamma({2\alpha+1\over 2})}.$ (28)
Replacing (27) and (28) in (26), yields
$\mathbb{E}(W)={2\Gamma(2\alpha)\over 2^{2\alpha}\alpha\Gamma^{2}(\alpha)}={\Gamma(2\alpha+1)\over\alpha^{2}2^{2\alpha}\Gamma^{2}(\alpha)},$
where in the last step we have used the well-known identity $\Gamma(x+1)=x\Gamma(x)$. By applying Legendre duplication formula (Abramowitz and Stegun,, 1972): $\Gamma(x)\Gamma\left(x+{1/2}\right)=2^{1-2x}\sqrt{\pi}\Gamma(2x),$ in the above identity, we get
$\mathbb{E}(W)={\Gamma(\alpha+{1\over 2})\over\sqrt{\pi}\alpha\Gamma(\alpha)},$
which corresponds to the Gini coefficient, $G$, formula for gamma distributions given in McDonald and Jensen, (1979).
6.2 The $n$th raw moment of $W$: case $r\in\mathbb{N}$ and $\lambda_{1}=\lambda_{2}$
By using Proposition A.4, with $\varepsilon=0$ and $\delta=\infty$, $\mathbb{E}(W^{n})$ can be written as
$\mathbb{E}(W^{n})$ $=n\int_{0}^{1}w^{n-1}[1-F_{W}(w)]{\rm d}w$
$=n\int_{0}^{1}w^{n-1}\left[1-\int_{1-w_{r}}^{w_{r}}f_{X\over X+Y}(u){\rm d}u\right]{\rm d}w$
$=4rn\int_{1/2}^{1}(2y-1)^{2rn-1}\left[1-\int_{1-y}^{y}f_{X\over X+Y}(u){\rm d}u\right]{\rm d}y,$
where in the second equality the formula (10) for $F_{W}$ was used, and in the third equality the change of variables $y=w_{r}={(1+w^{\!{}^{1/2r}})/2}$. Let $r\in\mathbb{N}$. By the Binomial Theorem, the above expression becomes
$=4rn\sum_{k=0}^{2rn-1}\binom{2rn-1}{k}2^{k}(-1)^{2rn-1-k}\int_{1/2}^{1}y^{k}\left[1-\int_{1-y}^{y}f_{X\over X+Y}(u){\rm d}u\right]{\rm d}y$
$=4rn\sum_{k=0}^{2rn-1}\binom{2rn-1}{k}2^{k}(-1)^{2rn-1-k}\xi(k),$
where $\xi(\cdot)$ is defined in (38). As $\lambda_{1}=\lambda_{2}$, by using Proposition A.5, the $n$th raw moment of $W$ can be written as
$\mathbb{E}(W^{n})=4rn\sum_{k=0}^{2rn-1}\binom{2rn-1}{k}{2^{k}(-1)^{2rn-1-k}\over k+1}\left\{{\mathrm{B}(\alpha_{1}+k+1,\alpha_{2})\over\mathrm{B}(\alpha_{1},\alpha_{2})}\left[1-I_{1/2}(\alpha_{1}+k+1,\alpha_{2})\right]\right.\\[5.69046pt] +\left.{\mathrm{B}(\alpha_{1},\alpha_{2}+k+1)\over\mathrm{B}(\alpha_{1},\alpha_{2})}\,I_{1/2}(\alpha_{1},\alpha_{2}+k+1)-\left({1\over 2}\right)^{k+1}\right\},$
provided $r\in\mathbb{N}$. In the above, $I_{x}(a,b)={\rm B}(x;a,b)/{\rm B}(a,b)$ denotes the regularized beta function, with ${\rm B}(x;a,b)$ being the incomplete beta function.
6.3 The 111st raw moment of $Z$: case $r=1/2$ and $\lambda_{2}/\lambda_{1}<2$
Under the condition $r=1/2$, in this subsection (more specifically, in (30) and Remark 6.3), we obtain analytical formulas for the 111st raw moment (expected value) of
$Z=1-{\sqrt{XY}\over{1\over 2}(X+Y)},$
where $X\sim\text{Gamma}(\alpha_{1},\lambda_{1})$ and $Y\sim\text{Gamma}(\alpha_{2},\lambda_{2})$ are independent.
Indeed, by using the identity (19) with $z=1/[(X+Y)/2]$, we have
$\mathbb{E}(Z)=1-\mathbb{E}\left[{\sqrt{XY}\over{1\over 2}(X+Y)}\right]$ $=1-\mathbb{E}\left[\sqrt{XY}\int_{0}^{\infty}\exp\left\{-\left({X+Y\over 2}\right)x\right\}{\rm d}x\right]$
$=1-\int_{0}^{\infty}\mathbb{E}\left[\sqrt{XY}\exp\left\{-\left({X+Y\over 2}\right)x\right\}\right]{\rm d}x,$
where we invoked Tonelli’s theorem (Folland,, 1999) in the last step to validate the change in integration order. By using the independence of $X$ and $Y$ the above expression becomes
$\mathbb{E}(Z)=1-\int_{0}^{\infty}\mathbb{E}\left[\sqrt{X}\exp\left\{-{x\over 2}\,X\right\}\right]\mathbb{E}\left[\sqrt{Y}\exp\left\{-{x\over 2}\,Y\right\}\right]{\rm d}x.$ (29)
Furthermore, as $X\sim\text{Gamma}(\alpha_{1},\lambda_{1})$, observe that
$\mathbb{E}\left[\sqrt{X}\exp\left\{-{x\over 2}\,X\right\}\right]={\lambda_{1}^{\alpha_{1}}\over\Gamma(\alpha_{1})}\,\int_{0}^{\infty}u^{\alpha_{1}+{1\over 2}-1}\exp\left\{-\left({x\over 2}+\lambda_{1}\right)u\right\}{\rm d}u={\lambda_{1}^{\alpha_{1}}\over\Gamma(\alpha_{1})}\,{\Gamma(\alpha_{1}+{1\over 2})\over\left({x\over 2}+\lambda_{1}\right)^{\alpha_{1}+{1\over 2}}}.$
Analogously, as $Y\sim\text{Gamma}(\alpha_{2},\lambda_{2})$,
$\mathbb{E}\left[\sqrt{Y}\exp\left\{-{x\over 2}\,Y\right\}\right]={\lambda_{2}^{\alpha_{2}}\over\Gamma(\alpha_{2})}\,{\Gamma(\alpha_{2}+{1\over 2})\over\left({x\over 2}+\lambda_{2}\right)^{\alpha_{2}+{1\over 2}}}.$
Replacing the last two relations in (29), we get
$\mathbb{E}(Z)$ $=1-{\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\Gamma(\alpha_{1}+{1\over 2})\Gamma(\alpha_{2}+{1\over 2})\over\Gamma(\alpha_{1})\Gamma(\alpha_{2})}\,\int_{0}^{\infty}{1\over\left({x\over 2}+\lambda_{1}\right)^{\alpha_{1}+{1\over 2}}\left({x\over 2}+\lambda_{2}\right)^{\alpha_{2}+{1\over 2}}}{\rm d}x$
$=1-{2\lambda_{2}^{\alpha_{2}}\Gamma(\alpha_{1}+{1\over 2})\Gamma(\alpha_{2}+{1\over 2})\over\lambda_{1}^{\alpha_{2}}(\alpha_{1}+\alpha_{2})\Gamma(\alpha_{1})\Gamma(\alpha_{2})}\,\,_{2}F_{1}\left(\alpha_{2}+{1\over 2},\alpha_{1}+\alpha_{2};\alpha_{1}+\alpha_{2}+1;1-{\lambda_{2}\over\lambda_{1}}\right),$ (30)
whenever $\lambda_{2}/\lambda_{1}<2$, where in the final equality, Proposition A.3 was applied.
Remark 6.3.
From (30), for $\lambda_{1}=\lambda_{2}$, we have
$\mathbb{E}(Z)=1-{2\Gamma(\alpha_{1}+{1\over 2})\Gamma(\alpha_{2}+{1\over 2})\over(\alpha_{1}+\alpha_{2})\Gamma(\alpha_{1})\Gamma(\alpha_{2})},$
because $\,{}_{2}F_{1}(a,b;c;0)=1$.
6.4 The $n$th raw moment of $Z$: case $r\in\mathbb{N}$ and $\lambda_{1}=\lambda_{2}$
Applying Proposition A.4, with $\varepsilon=0$ and $\delta=\infty$, $\mathbb{E}(Z^{n})$ becomes
$\mathbb{E}(Z^{n})$ $=n\int_{0}^{1}z^{n-1}[1-F_{Z}(z)]{\rm d}z$
$=n\int_{0}^{1}z^{n-1}\left[1-\int_{1-z_{r}}^{z_{r}}f_{X\over X+Y}(u){\rm d}u\right]{\rm d}z$
$=4rn\int_{1/2}^{1}[1-2^{r}(2y^{2}-y+1)^{r}]^{n-1}\left[1-\int_{1-y}^{y}f_{X\over X+Y}(u){\rm d}u\right]{\rm d}y,$
where in the second equality the formula (14) for $F_{Z}$ was used, and in the third equality the change of variables $y=z_{r}={[1+\sqrt{1-(1-z)^{1/r}}]/2}$. As $r\in\mathbb{N}$, by the multinomial Theorem the above expression is written as
$=4rn\sum_{k=0}^{n-1}\binom{n-1}{k}(-1)^{k}2^{rk}\sum_{i+j+l=rk}\binom{rk}{i,j,l}2^{i}(-1)^{j}\int_{1/2}^{1}y^{2i+j}\left[1-\int_{1-y}^{y}f_{X\over X+Y}(u){\rm d}u\right]{\rm d}y$
$=4rn\sum_{k=0}^{n-1}\binom{n-1}{k}(-1)^{k}2^{rk}\sum_{i+j+l=rk}\binom{rk}{i,j,l}2^{i}(-1)^{j}\xi(2i+j),$
where $\xi(\cdot)$ is as defined in (38) and
(${\displaystyle{n\choose k_{1},k_{2},\ldots,k_{m}}={\frac{n!}{k_{1}!\,k_{2}!\cdots k_{m}!}}}$
is a multinomial coefficient. As $\lambda_{1}=\lambda_{2}$, by using Proposition A.5, the $n$th raw moment of $Z$ can be written as
$\mathbb{E}(Z^{n})=4rn\sum_{k=0}^{n-1}\binom{n-1}{k}\sum_{i+j+l=rk}\binom{rk}{i,j,l}{2^{i+rk}(-1)^{j+k}\over 2i+j}\\[5.69046pt] \times\Bigg{\{}{\mathrm{B}(\alpha_{1}+2i+j+1,\alpha_{2})\over\mathrm{B}(\alpha_{1},\alpha_{2})}\left[1-I_{1/2}(\alpha_{1}+2i+j+1,\alpha_{2})\right]\\[5.69046pt] +{\mathrm{B}(\alpha_{1},\alpha_{2}+2i+j+1)\over\mathrm{B}(\alpha_{1},\alpha_{2})}\,I_{1/2}(\alpha_{1},\alpha_{2}+2i+j+1)-\left({1\over 2}\right)^{2i+j+1}\Bigg{\}},$
provided $r\in\mathbb{N}$. Here, $I_{x}(a,b)$ denotes the regularized beta function.
7 Maximum likelihood estimation
Let $\{W_{i}:i=1,\ldots,n\}$ and $\{Z_{i}:i=1,\ldots,n\}$ be univariate random samples of size $n$ from $W$ and $Z$ with density functions as given in (16) and (17), respectively, and let $w_{i}$ and $z_{i}$ be the sample observations of $W_{i}$ and $Z_{i}$, respectively. Then, the respective log-likelihood functions for $\bm{\theta}=(\alpha_{1},\alpha_{2},\lambda_{1},\lambda_{2},r)^{\top}$ are given by
$\ell(\bm{\theta};\bm{w})$ $=-\log(4)-\log(r)-{\alpha_{2}}\log(\lambda_{1})-{\alpha_{1}}\log(\lambda_{2})-\log(\text{B}(\alpha_{1},\alpha_{2}))$
$+\left({{1\over 2r}-1}\right)\sum_{i=1}^{n}\log(w_{i})+\sum_{i=1}^{n}\log(G(w_{r,i})),$ (31)
and
$\ell(\bm{\theta};\bm{z})$ $=-\log(4)-\log(r)-{\alpha_{2}}\log(\lambda_{1})-{\alpha_{1}}\log(\lambda_{2})-\log(\text{B}(\alpha_{1},\alpha_{2}))$
$+\left({{1\over r}-1}\right)\sum_{i=1}^{n}\log(1-z_{i})-{1\over 2}\sum_{i=1}^{n}\log\left({1-(1-z_{i})^{1\over r}}\right)+\sum_{i=1}^{n}\log(G(z_{r,i})),$ (32)
where
$w_{r,i}={1+w_{i}^{\!{}^{1\over 2r}}\over 2},\quad z_{r,i}={1+\sqrt{1-(1-z_{i})^{1\over r}}\over 2},\quad i=1,\ldots,n,$ (33)
and $G(\cdot)$ being as in (18). The necessary conditions for the occurrence of a maximum (or a minimum) of $\ell(\bm{\theta};\bm{w})$ and $\ell(\bm{\theta};\bm{z})$ are
${\partial\ell(\bm{\theta};\bm{w})\over\partial\alpha_{i}}=0,\quad{\partial\ell(\bm{\theta};\bm{w})\over\partial\lambda_{i}}=0,\quad{\partial\ell(\bm{\theta};\bm{w})\over\partial r}=0,\quad i=1,2,$ (34)
and
${\partial\ell(\bm{\theta};\bm{z})\over\partial\alpha_{i}}=0,\quad{\partial\ell(\bm{\theta};\bm{z})\over\partial\lambda_{i}}=0,\quad{\partial\ell(\bm{\theta};\bm{z})\over\partial r}=0,\quad i=1,2,$ (35)
respectively. The equations (34) and (35) are known as the likelihood equations. It is simple to observe that (for $i=1,2$)
${\partial\ell(\bm{\theta};\bm{w})\over\partial\alpha_{1}}=-\log(\lambda_{2})-[\psi(\alpha_{1})-\psi(\alpha_{1}+\alpha_{2})]+\sum_{i=1}^{n}{{\partial G(w_{r,i})\over\partial\alpha_{1}}\over G(w_{r,i})},$
${\partial\ell(\bm{\theta};\bm{w})\over\partial\alpha_{2}}=-\log(\lambda_{1})-[\psi(\alpha_{2})-\psi(\alpha_{1}+\alpha_{2})]+\sum_{i=1}^{n}{{\partial G(w_{r,i})\over\partial\alpha_{2}}\over G(w_{r,i})},$
${\partial\ell(\bm{\theta};\bm{w})\over\partial\lambda_{1}}=-{\alpha_{2}\over\lambda_{1}}+\sum_{i=1}^{n}{{\partial G(w_{r,i})\over\partial\lambda_{1}}\over G(w_{r,i})},$
${\partial\ell(\bm{\theta};\bm{w})\over\partial\lambda_{2}}=-{\alpha_{1}\over\lambda_{2}}+\sum_{i=1}^{n}{{\partial G(w_{r,i})\over\partial\lambda_{2}}\over G(w_{r,i})},$
${\partial\ell(\bm{\theta};\bm{w})\over\partial r}=-{1\over r}-{1\over 2r^{2}}\sum_{i=1}^{n}\log(w_{i})+\sum_{i=1}^{n}{G^{\prime}(w_{r,i})\over G(w_{r,i})}\,{\partial w_{r,i}\over\partial r},$
and
${\partial\ell(\bm{\theta};\bm{z})\over\partial\alpha_{1}}=-\log(\lambda_{2})-[\psi(\alpha_{1})-\psi(\alpha_{1}+\alpha_{2})]+\sum_{i=1}^{n}{{\partial G(z_{r,i})\over\partial\alpha_{1}}\over G(z_{r,i})},$
${\partial\ell(\bm{\theta};\bm{z})\over\partial\alpha_{2}}=-\log(\lambda_{1})-[\psi(\alpha_{2})-\psi(\alpha_{1}+\alpha_{2})]+\sum_{i=1}^{n}{{\partial G(z_{r,i})\over\partial\alpha_{2}}\over G(z_{r,i})},$
${\partial\ell(\bm{\theta};\bm{z})\over\partial\lambda_{1}}=-{\alpha_{2}\over\lambda_{1}}+\sum_{i=1}^{n}{{\partial G(z_{r,i})\over\partial\lambda_{1}}\over G(z_{r,i})},$
${\partial\ell(\bm{\theta};\bm{z})\over\partial\lambda_{2}}=-{\alpha_{1}\over\lambda_{2}}+\sum_{i=1}^{n}{{\partial G(z_{r,i})\over\partial\lambda_{2}}\over G(z_{r,i})},$
${\partial\ell(\bm{\theta};\bm{z})\over\partial r}=-{1\over r}-{1\over r^{2}}\sum_{i=1}^{n}\log(1-z_{i})-{1\over 2r^{2}}\sum_{i=1}^{n}{(1-z_{i})^{1\over r}\log(1-z_{i})\over 1-(1-z_{i})^{1\over r}}+\sum_{i=1}^{n}{G^{\prime}(z_{r,i})\over G(z_{r,i})}\,{\partial z_{r,i}\over\partial r},$
where $G^{\prime}(x)$ denotes the derivative of $G(x)$, given in (18), with respect to $x$, and ${\partial w_{r,i}/\partial r}$ and ${\partial z_{r,i}/\partial r}$ are the partial partial derivatives of $w_{r,i}$ and $z_{r,i}$ given in (33). In the above, $\psi(x)=\Gamma^{\prime}(x)/\Gamma(x)$ denotes the digamma function.
Any nontrivial root $\widehat{\bm{\theta}}=(\widehat{\alpha}_{1},\widehat{\alpha}_{2},\widehat{\lambda}_{1},\widehat{\lambda}_{2},\widehat{r})^{\top}$ of the above likelihood equations that provides the absolute maximum of the log-likelihood function is called an maximum likelihood (ML) estimate. Due to the complexity of the maximization problem, numerical optimization methods are required to obtain the ML estimate.
8 Conclusions
We have introduced two flexible families of unit‐interval distributions by applying non‐injective transformations to the ratio of independent gamma variates, obtaining closed‐form expressions for their densities and distribution functions. We have derived quantile representations, moments in special cases, and established a maximum likelihood estimation procedure for joint inference on $(\alpha_{1},\alpha_{2},\lambda_{1},\lambda_{2},r)$. As part of future research, it will be of interest to explore regression extensions, Bayesian methods, and applications to bounded‐response data.
Acknowledgements
The research was supported in part by CNPq and CAPES grants from the Brazilian government.
Disclosure statement
There are no conflicts of interest to disclose.
Appendix A Appendix
The proof of the following result is a direct consequence of the gamma distribution definition and is therefore omitted.
Proposition A.1.
If $X\sim\text{Gamma}(\alpha_{1},\lambda_{1})$ and $Y\sim\text{Gamma}(\alpha_{2},\lambda_{2})$, then, we have:
$\mathbb{E}\left[\exp(-Xx)\right]={\lambda_{1}^{\alpha_{1}}\over(x+\lambda_{1})^{\alpha_{1}}},\quad\mathbb{E}\left[X\exp(-Xx)\right]={\alpha_{1}\lambda_{1}^{\alpha_{1}}\over(x+\lambda_{1})^{\alpha_{1}+1}},$
and
$\mathbb{E}\left[\exp(-Yx)\right]={\lambda_{2}^{\alpha_{2}}\over(x+\lambda_{2})^{\alpha_{2}}},\quad\mathbb{E}\left[Y\exp(-Yx)\right]={\alpha_{2}\lambda_{2}^{\alpha_{2}}\over(x+\lambda_{2})^{\alpha_{2}+1}}.$
Proposition A.2.
If $X\sim\text{Gamma}(\alpha_{1},\lambda_{1})$ and $Y\sim\text{Gamma}(\alpha_{2},\lambda_{2})$, then, we have:
$\int_{0}^{\infty}\left[\int_{0}^{u}\exp(-vx){\rm d}F_{Y}(v)\right]$ $u\exp(-ux){\rm d}F_{X}(u)$
$={\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2}+1)}\,{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2}+1,1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}+1}},$
and
$\int_{0}^{\infty}\left[\int_{0}^{u}v\exp(-vx){\rm d}F_{Y}(v)\right]\exp(-ux){\rm d}F_{X}(u)$ $={\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\,{\,{}_{2}F_{1}\left(\alpha_{1}+\alpha_{2},1;\alpha_{2}+1;{x+\lambda_{2}\over 2x+\lambda_{1}+\lambda_{2}}\right)\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}$
$-{\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\text{B}(\alpha_{1},\alpha_{2})}\,{1\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}},$
[…]
${}_{2}F_{1}(a,b;c;z)=\sum_{n=0}^{\infty}{\frac{(a)_{n}(b)_{n}}{(c)_{n}}}{\frac{z^{n}}{n!}}.$ (36)
Here $(q)_{n}$ is the (rising) Pochhammer symbol, which is defined by:
${\displaystyle(q)_{n}={\begin{cases}1&n=0,\\ q(q+1)\cdots(q+n-1)&n>0.\end{cases}}}$
Proof.
It is simple to observe that
$\int_{0}^{u}\exp(-vx){\rm d}F_{Y}(v)={\lambda_{2}^{\alpha_{2}}\over\Gamma(\alpha_{2})(x+\lambda_{2})^{\alpha_{2}}}\,\gamma(\alpha_{2},(x+\lambda_{2})u)$
with $\gamma(s,x)=\int_{0}^{x}t^{s-1}\exp(-t){\rm d}t$ being the lower incomplete gamma function, and
$\int_{0}^{u}v\exp(-vx){\rm d}F_{Y}(v)$ $={\lambda_{2}^{\alpha_{2}}\over\Gamma(\alpha_{2})(x+\lambda_{2})^{\alpha_{2}+1}}\,\,\gamma(\alpha_{2}+1,(x+\lambda_{2})u)$
$={\lambda_{2}^{\alpha_{2}}\over\Gamma(\alpha_{2})(x+\lambda_{2})^{\alpha_{2}+1}}\,\left[\alpha_{2}\gamma(\alpha_{2},(x+\lambda_{2})u)-(x+\lambda_{2})^{\alpha_{2}}u^{\alpha_{2}}\exp\{-(x+\lambda_{2})u\}\right],$
where in the last line the well-known recurrence relation: $\gamma(s+1,x)=s\gamma(s,x)-x^{s}\exp(-x)$, have been used. Consequently,
$\int_{0}^{\infty}\left[\int_{0}^{u}\exp(-vx){\rm d}F_{Y}(v)\right]u\exp(-ux){\rm d}F_{X}(u)\\[5.69046pt] ={\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\Gamma(\alpha_{1})\Gamma(\alpha_{2})(x+\lambda_{2})^{\alpha_{2}}}\int_{0}^{\infty}u^{\alpha_{1}}\exp\{-(x+\lambda_{1})u\}\gamma(\alpha_{2},(x+\lambda_{2})u){\rm d}u$
and
$\int_{0}^{\infty}\Bigg{[}$ $\int_{0}^{u}v\exp(-vx){\rm d}F_{Y}(v)\Bigg{]}\exp(-ux){\rm d}F_{X}(u)$
$={\alpha_{2}\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\over\Gamma(\alpha_{1})\Gamma(\alpha_{2})(x+\lambda_{2})^{\alpha_{2}+1}}\int_{0}^{\infty}u^{\alpha_{1}-1}\exp\{-(x+\lambda_{1})u\}\gamma(\alpha_{2},(x+\lambda_{2})u){\rm d}u$
$-{\lambda_{1}^{\alpha_{1}}\lambda_{2}^{\alpha_{2}}\Gamma(\alpha_{1}+\alpha_{2})\over\Gamma(\alpha_{1})\Gamma(\alpha_{2})}\,{1\over(x+\lambda_{2})(2x+\lambda_{1}+\lambda_{2})^{\alpha_{1}+\alpha_{2}}}.$
Finally, by using the identity $\text{B}(a,b)=\Gamma(a)\Gamma(b)/\Gamma(a+b)$ and the relation (D’Aurizio,, 2016):
$\int_{0}^{\infty}x^{a-1}\exp(-sx)\gamma(b,\theta x){\rm d}x={\theta^{b}\Gamma(a+b)\over b(s+\theta)^{a+b}}\,_{2}F_{1}\left(a+b,1;b+1;{\theta\over s+\theta}\right),$
in the last two identities, the proof follows. ∎
Proposition A.3 (Technical result).
For $a,b,c,e>0$ such that $ae<2bc$, and for $p+q>1$, we have
$\int_{0}^{\infty}{1\over(ax+b)^{p}(cx+e)^{q}}{\rm d}x={a^{q-1}\over b^{p+q-1}c^{q}(p+q-1)}\,_{2}F_{1}\left(q,p+q-1;p+q;1-{ae\over bc}\right),$
where $\,{}_{2}F_{1}(a,b;c;z)$ is the hypergeometric function defined in (36).
Proof.
Making the change of variable $y=ax/b$, we have
$\int_{0}^{\infty}{1\over(ax+b)^{p}(cx+e)^{q}}{\rm d}x$ $={a^{q}\over ab^{p-1}c^{q}b^{q}}\int_{0}^{\infty}{1\over(y+1)^{p}({bc\over a}y+e)^{q}}{\rm d}y$
$={a^{q-1}\over b^{p+q-1}c^{q}}\int_{0}^{\infty}{1\over(y+1)^{p}(y+{ae\over bc})^{q}}{\rm d}y.$ (37)
Using the integral representation of hypergeometric functions:
$\,{}_{2}F_{1}(a,b;c;z)={\Gamma(c)\over\Gamma(b)\Gamma(c-b)}\int_{0}^{\infty}t^{-b+c-1}(t+1)^{a-c}(t-z+1)^{-a}{\rm d}t,\quad c>b>0,\ -1<z<1,$
with $a=q$, $b=p+q-1$, $c=p+q$ and $z=1-ae/cb$, expression in (A) becomes
$={a^{q-1}\over b^{p+q-1}c^{q}}{\Gamma(p+q-1)\over\Gamma(p+q)}\,_{2}F_{1}\left(q,p+q-1;p+q;1-{ae\over bc}\right),$
provided $ae<2bc$. Finally, invoking the well-known identity: $\Gamma(z+1)=z\Gamma(z)$, the proof follows. ∎
Proposition A.4 (Proposition 3.1 of Dutta et al., (2025)).
Let $X$ be a real-valued random variable and $p>0$ be a real number. For all $\varepsilon\geqslant 0$ and $\delta>0$, the following identity is satisfied:
$\mathbb{E}(X^{p}\mathds{1}_{\{\varepsilon<X<\delta\}})=\varepsilon^{p}\mathbb{P}(X>\varepsilon)-\delta^{p}\mathbb{P}(X>\delta)+p\int_{\varepsilon}^{\delta}u^{p-1}\mathbb{P}(X>u){\rm d}u.$
For simplicity, in the next proposition we adopt the following notation:
$\xi(p)\equiv\int_{1/2}^{1}y^{p}\left[1-\int_{1-y}^{y}f_{X\over X+Y}(u){\rm d}u\right]{\rm d}y.$ (38)
Proposition A.5.
If $X$ and $Y$ are independent gamma-distributed with parameter vectors $(\alpha_{1},\lambda_{1})^{\top}$ and $(\alpha_{2},\lambda_{2})^{\top}$, respectively, such that $\lambda_{1}=\lambda_{2}$, then
$\xi(p)={1\over p+1}\Bigg{\{}{\mathrm{B}(\alpha_{1}+p+1,\alpha_{2})\over\mathrm{B}(\alpha_{1},\alpha_{2})}\left[1-I_{1/2}(\alpha_{1}+p+1,\alpha_{2})\right]\\[5.69046pt] +{\mathrm{B}(\alpha_{1},\alpha_{2}+p+1)\over\mathrm{B}(\alpha_{1},\alpha_{2})}\,I_{1/2}(\alpha_{1},\alpha_{2}+p+1)-\left({1\over 2}\right)^{p+1}\Bigg{\}},$
where $p>\max\{-1,-1-\alpha_{1},-1-\alpha_{2}\}$. In the above, $I_{x}(a,b)={\rm B}(x;a,b)/{\rm B}(a,b)$ denotes the regularized beta function.
Proof.
Note that $\xi(p)$ in (38) can be written as
$\xi(p)=\int_{1/2}^{1}y^{p}\left[1-F_{X\over X+Y}(y)+F_{X\over X+Y}(1-y)\right]{\rm d}y.$
As $F_{X\over X+Y}(1-y)=1-F_{Y\over X+Y}(y)$, the above term simplifies to
$=\int_{1/2}^{1}y^{p}[1-F_{X\over X+Y}(y)]{\rm d}y+\int_{1/2}^{1}y^{p}[1-F_{Y\over X+Y}(y)]{\rm d}y.$
Applying Proposition A.4, the above expression becomes
$={1\over p+1}\left\{\mathbb{E}\left[\left({X\over X+Y}\right)^{p+1}\mathds{1}_{\{{1\over 2}<{X\over X+Y}<1\}}\right]-\left({1\over 2}\right)^{p+1}\mathbb{P}\left({X\over X+Y}>{1\over 2}\right)\right.$
$\hskip 170.71652pt\left.+\mathbb{E}\left[\left({Y\over X+Y}\right)^{p+1}\mathds{1}_{\{{1\over 2}<{Y\over X+Y}<1\}}\right]-\left({1\over 2}\right)^{p+1}\mathbb{P}\left({Y\over X+Y}>{1\over 2}\right)\right\}.$ (39)
Since $\lambda_{1}=\lambda_{2}$, then, it is clear that
$\mathbb{E}\left[\left({X\over X+Y}\right)^{p+1}\mathds{1}_{\{{1\over 2}<{X\over X+Y}<1\}}\right]$ $=\int_{1\over 2}^{1}\frac{z^{(\alpha_{1}+p+1)-1}(1-z)^{\alpha_{2}-1}}{\mathrm{B}(\alpha_{1},\alpha_{2})}{\rm d}z$
$={\mathrm{B}(\alpha_{1}+p+1,\alpha_{2})\over\mathrm{B}(\alpha_{1},\alpha_{2})}\int_{1\over 2}^{1}\frac{z^{(\alpha_{1}+p+1)-1}(1-z)^{\alpha_{2}-1}}{\mathrm{B}(\alpha_{1}+k+1,\alpha_{2})}{\rm d}z$
$={\mathrm{B}(\alpha_{1}+p+1,\alpha_{2})\over\mathrm{B}(\alpha_{1},\alpha_{2})}\left[1-I_{1/2}(\alpha_{1}+p+1,\alpha_{2})\right]$ (40)
and, analogously,
$\mathbb{E}\left[\left({Y\over X+Y}\right)^{p+1}\mathds{1}_{\{{1\over 2}<{Y\over X+Y}<1\}}\right]$ $={\mathrm{B}(\alpha_{1},\alpha_{2}+p+1)\over\mathrm{B}(\alpha_{1},\alpha_{2})}\left[1-I_{1/2}(\alpha_{2}+p+1,\alpha_{1})\right]$
$={\mathrm{B}(\alpha_{1},\alpha_{2}+p+1)\over\mathrm{B}(\alpha_{1},\alpha_{2})}\,I_{1/2}(\alpha_{1},\alpha_{2}+p+1),$ (41)
where in the last step the well-known identity $I_{x}(a,b)=1-I_{1-x}(b,a)$ was used.
Furthermore,
$\mathbb{P}\left({X\over X+Y}>{1\over 2}\right)=1-I_{1/2}(\alpha_{1},\alpha_{2}),\quad\mathbb{P}\left({Y\over X+Y}>{1\over 2}\right)=1-I_{1/2}(\alpha_{2},\alpha_{1})=I_{1/2}(\alpha_{1},\alpha_{2}).$ (42)
By plugging (A), (A) and (42) in (39), the proof follows. ∎